# EMAC+: Embodied Multimodal Agent for Collaborative Planning with VLM+LLM

### v1

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have transformed problem-solving in text-centric domains (Zhu2023MiniGPT4EV, ; Shinn2023ReflexionLA, ; Shridhar2020ALFWorldAT, ; Wang2023DescribeEP, ), facilitating complex planning, reasoning, and tool use via human directives. Nonetheless, implementing LLM-based agents for complex tasks Jiang2023VIMARM in real-world scenarios Huang2023VoxPoserC3 ; Mu2023EmbodiedGPTVP ; Sumers2023DistillingIV remains an open problem. Current multimodal agents typically treat LLMs as static planners, using them solely to generate symbolic instructions executed by predefined visual modules Driess2023PaLMEAE ; Dai2023InstructBLIPTG ; Li2023BLIP2BL . This approach neglects dynamic environmental feedback, hindering adaptability and limiting real-time refinements to action plans.
@@
Existing methodologies further exacerbate this disparity by regarding LLMs as static, frozen planners, thus separating their textual reasoning from environment dynamics. Models such as SayCan Ahn2022DoAI or Code-as-Policies Liang2022CodeAP exemplify this limitation, utilizing LLMs to generate symbolic action sequences without accounting for significant obstacles like obscured objects, kinematic constraints, or sensor noise. This static nature prevents LLMs from adapting based on interaction histories or updating their understanding of environment-specific constraints like torque limits or object occlusion patterns. Consequently, LLM-based agents frequently propose impractical actions due to a lack of situational awareness and real-time adaptability.
@@
One challenge is that LLM-based agents cannot be directly deployed in a visual world. While LLMs demonstrate impressive abilities in textual interactions, including planning, reflection, and reward shaping, their effectiveness drastically diminishes when directly applied to visual environments. Vision-Language Models (VLMs) have emerged to address this limitation by aligning LLM reasoning with visual inputs. However, existing VLMs predominantly rely on static image-text pair alignments during pretraining, which restricts their ability to understand dynamic visual environments effectively. Even advanced models such as GPT-4V 2023GPT4VisionSC struggle in embodied scenarios like ALFWorld, particularly in zero-shot settings Shridhar2020ALFWorldAT . In such cases, these models typically default to leveraging linguistic priors derived from detected objects rather than adapting plans based on task-specific visual dynamics and interactions.
@@
To bridge the gap of deploying LLM in the visual world while maintaining its strengths of both LLM and VLMs, in this paper, we introduce an Embodied Multi-modal Agent for Collaborative planning with LLM + VLM (EMAC+). EMAC+ allows continuous adaptation of high-level textual action plans based on real-time visual execution feedback. Unlike existing methods, this collaboration empowers the LLM to internalize environment-specific dynamics and improve its domain-specific comprehension, thus generating more accurate and feasible plans for complex robotic tasks. Our approach addresses critical limitations in previous static multimodal methods by providing dynamic, interactive updates to action planning through direct visual feedback. Moreover, with its understanding of physical dynamics, EMAC+ can also generate precise control actions in robotics tasks. While recent VLA models like PaLM-E Driess2023PaLMEAE also output low-level actions, our framework differs in two key ways: (1) A collaborative scheme for LLM + VLM: EMAC+ first uses the prior knowledge from an LLM expert to predict the action sequences. The VLM agent will execute the planned actions to interact with the visual environment, and provide feedback for the LLM expert, including the knowledge of the history trajectories and the environment dynamics. (2) Domain-aware planning: Due to VLM feedback, our LLM expert will possess both contextual knowledge and practical experience of the agent’s execution. In contrast to PaLM-E’s open-loop execution, our LLM systematically refines its plans based on feedback from the VLM. This closed-loop synergy ensures that both models change simultaneously, with the LLM changing to suggest actions that work with the VLM’s control abilities.
@@
We validate EMAC+ in two benchmarks: ALFWorld Shridhar2020ALFWorldAT , which includes numerous tasks in both visual and textual environments, and RT-1 Brohan2022RT1RT , a real-world robotic platform for manipulation. Experiments demonstrate that EMAC+ achieves better performance in both benchmarks with a higher learning efficiency. Meanwhile, EMAC+ demonstrates enhanced generalization capability for out-of-distribution tasks, and exhibits more consistent performance amid environment disturbances. These results highlight three key contributions: (1) a bidirectional training paradigm that harmonizes symbolic reasoning with vision-driven control, enabling the LLM to learn environment dynamics through the VLM’s feedback; (2) an embodied agent that directly generates low-level actions from pixels while retaining the interpretability of language-guided planning; and (3) validation of a co-adaptive LLM-VLM framework in both simulated and real physical environments. Our research propels the boundaries of LLMs as adaptive controllers, facilitating the development of resilient, generalizable agents that acquire knowledge through visual interactions while preserving the natural language-driven reasoning.
@@ 2 Related Works
A key hurdle for LLM/VLM agents is their inability to generalize beyond training distributions. For example, CLIPort Shridhar2021CLIPortWA demonstrates strong performance on tabletop tasks but fails with novel object configurations or lighting conditions. Similarly, VIMA Jiang2023VIMARM highlights that VLMs trained on static image-text pairs struggle to adapt to dynamic, partially observable RL environments. Recent work by  lin2025sim identifies a “sim-to-real perception gap” where agents overfit to simulated visuals and fail under real-world sensory noise. LLM-generated plans (e.g., “pick up the tool”) often lack precision for low-level control, leading to execution failures. Efforts like Distilled Behavior Cloning Liang2024MonitoringAC attempt to mitigate these issues by distilling LLM knowledge into RL policies, but they sacrifice the flexibility of direct language grounding. Due to the nature of language, the textual descriptions of pixel observations are prone to losing important details for robotic control. Unlike these methods, EMAC+ integrates real-time visual feedback into a collaborative loop between LLM and VLM, allowing continuous refinement of plans based on dynamic observations. This interactive process enables EMAC+ to adapt flexibly to novel scenarios and environmental perturbations, thereby significantly mitigating the generalization issues traditional static-alignment models face. Consequently, EMAC+ can effectively translate high-level textual instructions into precise, executable actions suitable for robust robotic control tasks.
@@
Aligning LLMs and VLMs requires bridging semantic and embodiment gaps. Semantically, LLMs reason in abstract language, while VLMs map pixels to text, creating a mismatch in representational granularity. For instance, LLaVA Liu2023VisualIT aligns LLMs with VLMs for visual QA but lacks mechanisms to translate answers into actions. Architectures like PaLM-E Driess2023PaLMEAE embed both modalities into a shared latent space but struggle with real-time policy updates during RL training. More specifically, the reward design in alignment is challenging, i.e., LLM-generated rewards (e.g., the correctness of step-by-step plans) often misalign with environment-specific success signals Wang2023DescribeEP . Then, VLMs process static frames, while LLMs generate sequence-based plans, leading to incoherent action sequences in long-horizon tasks. On the other hand, jointly fine-tuning LLMs and VLMs risks catastrophic forgetting of pre-trained knowledge. Moreover, VLMs Li2023MultimodalFM enable plans to be conditioned on visual inputs that are converted into linguistic descriptions or token embeddings aligned with LLMs Ahn2022DoAI ; Gao2023PhysicallyGV ; Huang2023VoxPoserC3 ; Helmert2006TheFD . Nevertheless, current foundation models are typically pretrained on static text or text-image datasets Driess2023PaLMEAE ; Mu2023EmbodiedGPTVP ; Yang2023OctopusEV ; Zhong2024PolicyIU , which may hinder their ability to adapt to the world’s changes. Due to the nature of language, the textual descriptions of pixel observations are prone to losing important details for robotic control. EMAC+ mitigates the issue of imprudent planning activities resulting from absent critical knowledge by analyzing both short-term and long-term historical data, encompassing the historical trajectories of the interaction between the VLM and its environment. Simultaneously, by analyzing failure paths, the LLM agent will avoid the forgetting problems and actively comprehend the environmental physical knowledge, enhancing the planning quality.
@@ 3 Methodology
Fig. 1 illustrates the main idea of our multi-modal agent, who trained by the interactive knowledge between LLM and VLM. Our agent is built on a modularized VLM, which can take pixel observations as input and generate low-level control actions. We follow  Yang2023EmbodiedMA to construct a parallel text world for providing EMAC+ with step guidance to achieve this and overcome the associated challenges. Moreover, we design an interactive training scheme for the LLM expert to allow it to have a clear understanding of the gap between the text world and the real visual environment.
@@ 3.1 Multi-Modal Agent
In visual environments, the multi-modal agent $\pi_{\theta}$ is designed to process a task instruction $x_{task}$, and the pixel observation $s_{v}^{t}$ at each interaction step $t$. $\pi_{\theta}$ is trained to generate a sequence of real control actions $\left\{x_{a}^{t}\sim\pi_{\theta}(\cdot|x_{task},s_{v}^{t})\right\}$ to control the agent/robot to complete the task efficiently. To achieve this, inspired by recent works (Dai2023InstructBLIPTG, ; Li2023MultimodalFM, ; Li2023BLIP2BL, ; Liu2023VisualIT, ; Yang2023EmbodiedMA, ; Zhu2023MiniGPT4EV, ), we design the VLM-based agent of five components: (1) a ViT to encode $s_{v}$ into visual embeddings, (2) a querying transformer (Q-Former) tailored to extract the most relevant visual features via the cross-attention between the visual embeddings and query tokens, (3) a linear projection layer to align visual features to text embeddings, (4) an LLM decoder taking the concatenation of the instruction tokens and the output of the linear projection layer to autoregressively generate the action $x_{a}$, (5) a pre-trained action dictionary mapping the textual actions to control actions. In our model, we adopt the pre-trained ViT, Q-former, and LLM from InstructBLIP (Dai2023InstructBLIPTG, ) and keep them frozen at the finetuning stage. Note that we only update the linear projection layer as shown in Fig. 1. To enable our multi-modal agent to work in different tasks, e.g., manipulation tasks, which require the model to generate control actions. Given the previous success of PaLM-E (Driess2023PaLMEAE, ), it is reasonable to assume that we have access to policies that can perform low-level skills from some (small) vocabulary as long as the multi-modal agent consists of a sequence of such skills. Thus, we adopt an action mapping dictionary that contains low-level skills (e.g., “rotate the arm with $90^{\circ}$”), which maps such skills to control actions.
@@
Nonetheless, there are persistent issues that require our attention, as outlined in Sec. 1. While the pre-trained VLM (Dai2023InstructBLIPTG, ; Li2023MultimodalFM, ; Li2023BLIP2BL, ) is functional, it is suboptimal, as the pre-training process restricts the model to focus solely on the static alignment of image-text pairs, and the trained agent cannot reason about the dynamics of the environment. For example, even GPT-4V, a SOTA VLM, fails to accomplish tasks in an embodied visual environment that contains complex dynamics and object relationships (e.g., ALFWorld (Li2023MultimodalFM, ) in our experiments). More specifically, the primary reason for the failure of VLM agents is their reliance on the logical relationships of currently observed objects to create action instructions while neglecting the task instruction and the underlying dynamics of the environment. The strength of LLM lies in its ability to identify similarities and commonalities despite variations in details across multiple environments. Furthermore, finetuning a pre-trained VLM on a pre-collected dataset is still suboptimal due to the variability of environments and tasks (Brohan2023RT2VM, ), the absence of large-scale expert annotations (Padalkar2023OpenXR, ), and the challenges arising from distribution shift issues (Levine2020OfflineRL, ).
@@ 3.2 Text descriptions on visual observations
A core challenge in aligning vision-language models (VLMs) with large language models (LLMs) for manipulation tasks is bridging the gap between pixel-level observations and the symbolic reasoning required by LLMs. A common way to overcome this gap is to utilize PDDL (Planning Domain Definition Language) to convert raw visual observations into structured textual states Yang2023EmbodiedMA ; Shridhar2020ALFWorldAT . This enables the LLM to interpret the environment symbolically while retaining actionable semantics for downstream planning. The translation module integrates key attributes, such as Observed Objects, observed relations, inventory, and locations. Nevertheless, the text derived from pixel observation can solely furnish the LLM with abstract and ambiguous descriptions, merely encapsulating the fundamental information in the environment but failing to convey the physical specifics of the particular environment, task, and comprehensive action descriptions. While textual descriptions suffice for certain jobs (such as alfworld Shridhar2020ALFWorldAT and epic-kitchen Damen2018epickitchen ), they are inadequate for robotic control tasks that necessitate greater precision. Therefore, in EMAC+, we set up a retrospective feedback for the LLM agent to build a closed-loop interaction between the LLM agent and the VLM agent. Within this loop, the LLM agent not only has the short-term memory of planned actions, but it will also have the long-term memory of the historical trajectories of the agent in the environment for re-planning, which allows the LLM agent to understand the dynamics of the visual world and change the planned action sequences accordingly.
@@ 3.3 Build Collaborative LLM and VLM
Given the parallel text world translated from the visual world, we aim to train a VLM agent $\pi_{\theta}$ to complete tasks in the visual world. To align the prior knowledge provided by the LLM expert with the VLM, inspired by  Yang2023EmbodiedMA , we firstly formulate it as an imitation learning, and the objective can be written as:
@@
in which $\mathcal{L}$ is dependent on the specific tasks. In EMAC+, we use DPO (Rafailov2023DirectPO, ) loss as its proven superior performance in the cross-entropy on aligning models with expert preferences within the discrete language space. Thus, we rewrite Eq. 1 as:
@@
Given a pre-trained VLM agent learned from an expert, we now build a closed-loop collaboration between VLM and LLM. Given a task instruction $x_{task}$ and the textual observation $s_{l}$ translated by PDDL, the LLM agent first generates a high-level textual action sequence $x_{a,1:N}$ that is expected to complete $x_{task}$ step-by-step. LLM also infers low-level execution $b$ (one step action) given $=x_{a,1:N}$ and $s_{l}$. VLM agent executes the control actions to interact with the environment based on $b$. After the execution, the environment returns a new pixel observation $s_{v}$. LLM agent will decide whether it needs to modify $x_{a,1:N}$ to $x^{*}_{a,1:N}$ based on $s_{v}$ and the corresponding text state $s_{l}$. Thus, we collect the trajectories $\mathcal{D}=\left\{(s_{t},g,x_{a,1:N},x^{*}_{a,1:N}\right\}$. We fine-tune the LLM with LoRA (Hu2021LoRALA, ) for parameter-efficient updates, retaining the LLM’s general knowledge while adapting to environment dynamics as:

### v2

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have transformed problem-solving in text-centric domains (Zhu et al., 2023; Shinn et al., 2023; Shridhar et al., 2020; Wang et al., 2023), facilitating complex planning, reasoning, and tool use via human directives. Nonetheless, implementing LLM-based agents for complex tasks Jiang et al. (2023) in real-world scenarios Huang et al. (2023); Mu et al. (2023); Sumers et al. (2023) remains an open problem. Current multimodal agents typically treat LLMs as static planners, using them solely to generate symbolic instructions executed by predefined visual modules Driess et al. (2023); Dai et al. (2023); Li et al. (2023b). This approach neglects dynamic environmental feedback, hindering adaptability and limiting real-time refinements to action plans.
@@
Existing methodologies further exacerbate this disparity by regarding LLMs as static, frozen planners, thus separating their textual reasoning from environment dynamics. Models such as SayCan Ahn et al. (2022) or Code-as-Policies Liang et al. (2022) exemplify this limitation, utilizing LLMs to generate symbolic action sequences without accounting for significant obstacles like obscured objects, kinematic constraints, or sensor noise. This static nature prevents LLMs from adapting based on interaction histories or updating their understanding of environment-specific constraints like torque limits or object occlusion patterns. Consequently, LLM-based agents frequently propose impractical actions due to a lack of situational awareness and real-time adaptability.
@@
One challenge is that LLM-based agents cannot be directly deployed in a visual world. While LLMs demonstrate impressive abilities in textual interactions, including planning, reflection, and reward shaping, their effectiveness drastically diminishes when directly applied to visual environments. Vision-Language Models (VLMs) have emerged to address this limitation by aligning LLM reasoning with visual inputs. However, existing VLMs predominantly rely on static image-text pair alignments during pretraining, which restricts their ability to understand dynamic visual environments effectively. Even advanced models such as GPT-4V 202 (2023) struggle in embodied scenarios like ALFWorld, particularly in zero-shot settings Shridhar et al. (2020). In such cases, these models typically default to leveraging linguistic priors derived from detected objects rather than adapting plans based on task-specific visual dynamics and interactions.
@@
To bridge the gap of deploying LLM in the visual world while maintaining its strengths of both LLM and VLMs, in this paper, we introduce an Embodied Multi-modal Agent for Collaborative planning with LLM + VLM (EMAC+). EMAC+ allows continuous adaptation of high-level textual action plans based on real-time visual execution feedback. Unlike existing methods, this collaboration empowers the LLM to internalize environment-specific dynamics and improve its domain-specific comprehension, thus generating more accurate and feasible plans for complex robotic tasks. Our approach addresses critical limitations in previous static multimodal methods by providing dynamic, interactive updates to action planning through direct visual feedback. Moreover, with its understanding of physical dynamics, EMAC+ can also generate precise control actions in robotics tasks. While recent VLA models like PaLM-E Driess et al. (2023) also output low-level actions, our framework differs in two key ways: (1) A collaborative scheme for LLM + VLM: EMAC+ first uses the prior knowledge from an LLM expert to predict the action sequences. The VLM agent will execute the planned actions to interact with the visual environment, and provide feedback for the LLM expert, including the knowledge of the history trajectories and the environment dynamics. (2) Domain-aware planning: Due to VLM feedback, our LLM expert will possess both contextual knowledge and practical experience of the agent’s execution. In contrast to PaLM-E’s open-loop execution, our LLM systematically refines its plans based on feedback from the VLM. This closed-loop synergy ensures that both models change simultaneously, with the LLM changing to suggest actions that work with the VLM’s control abilities.
@@
We validate EMAC+ in two benchmarks: ALFWorld Shridhar et al. (2020), which includes numerous tasks in both visual and textual environments, and RT-1 Brohan et al. (2022), a real-world robotic platform for manipulation. Experiments demonstrate that EMAC+ achieves better performance in both benchmarks with a higher learning efficiency. Meanwhile, EMAC+ demonstrates enhanced generalization capability for out-of-distribution tasks, and exhibits more consistent performance amid environment disturbances. These results highlight three key contributions: (1) a bidirectional training paradigm that harmonizes symbolic reasoning with vision-driven control, enabling the LLM to learn environment dynamics through the VLM’s feedback; (2) an embodied agent that directly generates low-level actions from pixels while retaining the interpretability of language-guided planning; and (3) validation of a co-adaptive LLM-VLM framework in both simulated and real physical environments. Our research propels the boundaries of LLMs as adaptive controllers, facilitating the development of resilient, generalizable agents that acquire knowledge through visual interactions while preserving the natural language-driven reasoning.
@@ 2 Related Works
A key hurdle for LLM/VLM agents is their inability to generalize beyond training distributions. For example, CLIPort Shridhar et al. (2021) demonstrates strong performance on tabletop tasks but fails with novel object configurations or lighting conditions. Similarly, VIMA Jiang et al. (2023) highlights that VLMs trained on static image-text pairs struggle to adapt to dynamic, partially observable RL environments. Recent work by  Lin et al. (2025) identifies a “sim-to-real perception gap” where agents overfit to simulated visuals and fail under real-world sensory noise. LLM-generated plans (e.g., “pick up the tool”) often lack precision for low-level control, leading to execution failures. Efforts like Distilled Behavior Cloning Liang et al. (2024) attempt to mitigate these issues by distilling LLM knowledge into RL policies, but they sacrifice the flexibility of direct language grounding. Due to the nature of language, the textual descriptions of pixel observations are prone to losing important details for robotic control. Unlike these methods, EMAC+ integrates real-time visual feedback into a collaborative loop between LLM and VLM, allowing continuous refinement of plans based on dynamic observations. This interactive process enables EMAC+ to adapt flexibly to novel scenarios and environmental perturbations, thereby significantly mitigating the generalization issues traditional static-alignment models face. Consequently, EMAC+ can effectively translate high-level textual instructions into precise, executable actions suitable for robust robotic control tasks.
@@
Aligning LLMs and VLMs requires bridging semantic and embodiment gaps. Semantically, LLMs reason in abstract language, while VLMs map pixels to text, creating a mismatch in representational granularity. For instance, LLaVA Liu et al. (2023) aligns LLMs with VLMs for visual QA but lacks mechanisms to translate answers into actions. Architectures like PaLM-E Driess et al. (2023) embed both modalities into a shared latent space but struggle with real-time policy updates during RL training. More specifically, the reward design in alignment is challenging, i.e., LLM-generated rewards (e.g., the correctness of step-by-step plans) often misalign with environment-specific success signals Wang et al. (2023). Then, VLMs process static frames, while LLMs generate sequence-based plans, leading to incoherent action sequences in long-horizon tasks. On the other hand, jointly fine-tuning LLMs and VLMs risks catastrophic forgetting of pre-trained knowledge. Moreover, VLMs Li et al. (2023a) enable plans to be conditioned on visual inputs that are converted into linguistic descriptions or token embeddings aligned with LLMs Ahn et al. (2022); Gao et al. (2023a); Huang et al. (2023); Helmert (2006). Nevertheless, current foundation models are typically pretrained on static text or text-image datasets Driess et al. (2023); Mu et al. (2023); Yang et al. (2023a); Zhong et al. (2024), which may hinder their ability to adapt to the world’s changes. Due to the nature of language, the textual descriptions of pixel observations are prone to losing important details for robotic control. EMAC+ mitigates the issue of imprudent planning activities resulting from absent critical knowledge by analyzing both short-term and long-term historical data, encompassing the historical trajectories of the interaction between the VLM and its environment. Simultaneously, by analyzing failure paths, the LLM agent will avoid the forgetting problems and actively comprehend the environmental physical knowledge, enhancing the planning quality.
@@ 3 Methodology
Fig. 1 illustrates the main idea of our multi-modal agent, who trained by the interactive knowledge between LLM and VLM. Our agent is built on a modularized VLM, which can take pixel observations as input and generate low-level control actions. We follow  Yang et al. (2023b) to construct a parallel text world for providing EMAC+ with step guidance to achieve this and overcome the associated challenges. Moreover, we design an interactive training scheme for the LLM expert to allow it to have a clear understanding of the gap between the text world and the real visual environment.
@@ 3.1 Multi-Modal Agent
In visual environments, the multi-modal agent $\pi_{\theta}$ is designed to process a task instruction $x_{task}$, and the pixel observation $s_{v}^{t}$ at each interaction step $t$. $\pi_{\theta}$ is trained to generate a sequence of real control actions $\left\{x_{a}^{t}\sim\pi_{\theta}(\cdot|x_{task},s_{v}^{t})\right\}$ to control the agent/robot to complete the task efficiently. To achieve this, inspired by recent works (Dai et al., 2023; Li et al., 2023a; b; Liu et al., 2023; Yang et al., 2023b; Zhu et al., 2023), we design the VLM-based agent of five components: (1) a ViT to encode $s_{v}$ into visual embeddings, (2) a querying transformer (Q-Former) tailored to extract the most relevant visual features via the cross-attention between the visual embeddings and query tokens, (3) a linear projection layer to align visual features to text embeddings, (4) an LLM decoder taking the concatenation of the instruction tokens and the output of the linear projection layer to autoregressively generate the action $x_{a}$, (5) a pre-trained action dictionary mapping the textual actions to control actions. In our model, we adopt the pre-trained ViT, Q-former, and LLM from InstructBLIP (Dai et al., 2023) and keep them frozen at the finetuning stage. Note that we only update the linear projection layer as shown in Fig. 1. To enable our multi-modal agent to work in different tasks, e.g., manipulation tasks, which require the model to generate control actions. Given the previous success of PaLM-E (Driess et al., 2023), it is reasonable to assume that we have access to policies that can perform low-level skills from some (small) vocabulary as long as the multi-modal agent consists of a sequence of such skills. Thus, we adopt an action mapping dictionary that contains low-level skills (e.g., “rotate the arm with $90^{\circ}$”), which maps such skills to control actions.
@@
Nonetheless, there are persistent issues that require our attention, as outlined in Sec. 1. While the pre-trained VLM (Dai et al., 2023; Li et al., 2023a; b) is functional, it is suboptimal, as the pre-training process restricts the model to focus solely on the static alignment of image-text pairs, and the trained agent cannot reason about the dynamics of the environment. For example, even GPT-4V, a SOTA VLM, fails to accomplish tasks in an embodied visual environment that contains complex dynamics and object relationships (e.g., ALFWorld (Li et al., 2023a) in our experiments). More specifically, the primary reason for the failure of VLM agents is their reliance on the logical relationships of currently observed objects to create action instructions while neglecting the task instruction and the underlying dynamics of the environment. The strength of LLM lies in its ability to identify similarities and commonalities despite variations in details across multiple environments. Furthermore, finetuning a pre-trained VLM on a pre-collected dataset is still suboptimal due to the variability of environments and tasks (Brohan et al., 2023), the absence of large-scale expert annotations (Padalkar et al., 2023), and the challenges arising from distribution shift issues (Levine et al., 2020).
@@ 3.2 Text descriptions on visual observations
A core challenge in aligning vision-language models (VLMs) with large language models (LLMs) for manipulation tasks is bridging the gap between pixel-level observations and the symbolic reasoning required by LLMs. A common way to overcome this gap is to utilize PDDL (Planning Domain Definition Language) to convert raw visual observations into structured textual states Yang et al. (2023b); Shridhar et al. (2020). This enables the LLM to interpret the environment symbolically while retaining actionable semantics for downstream planning. The translation module integrates key attributes, such as Observed Objects, observed relations, inventory, and locations. Nevertheless, the text derived from pixel observation can solely furnish the LLM with abstract and ambiguous descriptions, merely encapsulating the fundamental information in the environment but failing to convey the physical specifics of the particular environment, task, and comprehensive action descriptions. While textual descriptions suffice for certain jobs (such as alfworld Shridhar et al. (2020) and epic-kitchen Damen et al. (2018)), they are inadequate for robotic control tasks that necessitate greater precision. Therefore, in EMAC+, we set up a retrospective feedback for the LLM agent to build a closed-loop interaction between the LLM agent and the VLM agent. Within this loop, the LLM agent not only has the short-term memory of planned actions, but it will also have the long-term memory of the historical trajectories of the agent in the environment for re-planning, which allows the LLM agent to understand the dynamics of the visual world and change the planned action sequences accordingly.
@@ 3.3 Build Collaborative LLM and VLM
Given the parallel text world translated from the visual world, we aim to train a VLM agent $\pi_{\theta}$ to complete tasks in the visual world. To align the prior knowledge provided by the LLM expert with the VLM, inspired by  Yang et al. (2023b), we firstly formulate it as an imitation learning, and the objective can be written as:
@@
in which $\mathcal{L}$ is dependent on the specific tasks. In EMAC+, we use DPO (Rafailov et al., 2023) loss as its proven superior performance in the cross-entropy on aligning models with expert preferences within the discrete language space. Thus, we rewrite Eq. 1 as:
@@
Given a pre-trained VLM agent learned from an expert, we now build a closed-loop collaboration between VLM and LLM. Given a task instruction $x_{task}$ and the textual observation $s_{l}$ translated by PDDL, the LLM agent first generates a high-level textual action sequence $x_{a,1:N}$ that is expected to complete $x_{task}$ step-by-step. LLM also infers low-level execution $b$ (one step action) given $x_{a,1:N}$ and $s_{l}$. VLM agent executes the control actions to interact with the environment based on $b$. After the execution, the environment returns a new pixel observation $s_{v}$. LLM agent will decide whether it needs to modify $x_{a,1:N}$ to $x^{*}_{a,1:N}$ based on $s_{v}$ and the corresponding text state $s_{l}$. Thus, we collect the trajectories $\mathcal{D}=\left\{(s_{t},g,x_{a,1:N},x^{*}_{a,1:N}\right\}$. We fine-tune the LLM with LoRA (Hu et al., 2021) for parameter-efficient updates, retaining the LLM’s general knowledge while adapting to environment dynamics as: